OpenStudy (keannathomas1):

Write an equation of a line perpendicular to line QR in slope intercept form that contains point (5,6)

OpenStudy (thereisnousername):

I assume the equation is y = (-1/2)x + 1. This equation is in the standard slope-intercept form: y = mx + b Where m is the slope of the line and b is the y-intercept. In your equation, m = -1/2 and b= 1. The slope of a line perpendicular to line QR will have a slope of -(1/m) = -1/(-1/2) = 2. So the equation of the line perpendicular to line QR is: y = 2x + b To find b, insert the point (5,6) that the line contains as defined in the problem statement and solve for b: y = 2x + b 6 = 2(5) + b [Insert x=5, y=6] 6 = 10 + b -4 = b [Subtract 10 from both sides So the final equation for the line perpendicular to line QR is: y = 2x - 4

OpenStudy (thereisnousername):

okay?

OpenStudy (keannathomas1):

Yes thank you